# Fabien Despinoy

### v2

@@ 3.2 Stage 1: Object detection & masks generation
Stage 1.1: Querying SAM with CLIP
High-resolution feature extraction We (over-)sample each image into a high resolution, divide it into $C$ crops in a sliding window fashion, process each crop separately with CLIP and rearrange the patches returned from CLIP (section 3.1) for all crops to construct the features for the full image. This guarantees precise and high quality feature maps for each image. Besides, for each crop $c$, we extract a classification token $\mathbf{cls}_{c}$.
Defining the set of concerned categories. To detect classes present in an image our method computes the similarity between the classification token for each crop to a set of text features corresponding to descriptions extracted from vocabulary $V$ (section 3.1). All possible labels in a given dataset are considered as the vocabulary (e.g. a set of 171 classes for COCO-Stuff).
Object presence detection. Each crop $c$ is classified with an object category when the object’s text feature $\textbf{t}_{o}$ has the highest similarity to the classification token $\mathbf{cls}_{c}$ among all descriptions extracted from vocabulary $V$. An object is considered as present in the image if it has been assigned to more than a predefined number of crops (set to 1 in our experiments).
Pseudo mask generation. We compute a similarity matrix between the full image feature map and the text features for the detected categories (Fig. 2). For each category we select 5 patches with the highest similarity as query points. We feed these points along with the original image to SAM and select the mask with the highest confidence.
Stage 1.2: SAM masks classification
Stage 1.1 may ignore small objects or generate partial masks for an image (Fig. 3) due to sub-optimal query points. Stage 1.2 performs a complementary pseudo label generation using CLIP and SAM to further boost the performance. We refer to the supplementary for comparison of the performance for stage 1.1 and stage 1.2.
@@
Mask labelling In order to classify each generated mask by SAM, we compute CLIP’s mean feature embedding corresponding to the area covered by the mask and compute its similarity to the text features of the detected categories in the image. The class with the highest similarity is selected as the pseudo label for the corresponding mask.
@@ 3.3 Stage 2: Lightweight semantic segmentation
Stage 1 extracts segmentation masks from a dataset with a predefined vocabulary. However, the generated masks might be noisy. Additionally, it involves querying two foundation models and might render inefficient when constrained with low compute. To overcome this, we propose to use annotations generated in Stage 1 to align any off-the-shelf pretrained vision encoder with text semantics and with no human supervision.
@@
Alignment module. We consider the generated masks in 3.2 as pseudo labels for training a simple alignment module that maps the image patch features to the text embedding. Such a mapping would ground any pretrained vision encoder with language for dense prediction tasks. Since supervised models may show a bias towards some object categories, we focus on self-supervised pretrained models. Namely, we consider DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.], a recent model trained in a fully self-supervised fashion with no text alignment. To overcome the noisy pseudo annotations we rely on 1) frozen pretrained text features as anchors, 2) already discriminative image patches features 3) a loss function that is robust to noise.
[…]
@@
Training the alignment module. The pretrained image encoder remains frozen and we optimize a small alignment module $\mathcal{M}$ to map the patch representations $\{\mathbf{x}^{p}_{i}\}$ to the CLIP text embedding space $z^{p}_{i}=\mathcal{M}(\mathbf{x}^{p}_{i})\in\mathbb{R}^{D}$. Note that the method is agnostic to the specific encoder used. As we strive for simplicity, we design our alignment module as one transformer block with multi-head self-attention layer. The self-attention layer allows each patch to attend to other patches in the image. Since we aim for an open-set semantic segmentation, cross-attention with text features [Alayrac et al.(2022)Alayrac, Donahue, Luc, Miech, Barr, Hasson, Lenc, Mensch, Millican, Reynolds, et al., Zou et al.(2023)Zou, Dou, Yang, Gan, Li, Li, Dai, Behl, Wang, Yuan, et al.] is not used as it would require a joint processing of the image features and a closed set of text features at test time.
@@
For notation clarity, we drop the image index and consider only an across-batch patch index $i;i\in\{1,\dots,N*B\}$ where N is the number of patches in an image and $B$ is the batch size. CLIP [He et al.(2023)He, Jamonnak, Gou, and Ren] contrasts the similarity of text features with image class tokens using a cross entropy loss, where a one to one correspondence exists between an image class token and its text features. In our case, we have many image features $\mathbf{z}_{i}=\mathcal{M}(\mathbf{x}_{i})$ extracted from patches of many images and few text features corresponding to detected categories in $\mathbf{D}$. Upon early experiments with CLIP loss, we found it not scalable to image patches and exhibiting poor convergence.
@@
Inspired by supervised contrastive loss SupCon [Khosla et al.(2020)Khosla, Teterwak, Wang, Sarna, Tian, Isola, Maschinot, Liu, and Krishnan] that operates on positive and negative pairs, we construct two types of pairs: pairs of image patches (patch-patch pairs: $<\mathbf{z}_{i},\mathbf{z}_{j}>$) and pairs of image patches and text features (patch-text pairs: $<\mathbf{z}_{i},\mathbf{t}_{k}>$). A patch-patch pair $<\mathbf{z}_{i},\mathbf{z}_{j}>$ is considered positive if the patches belong to the same category class $y_{i}=y_{j}$ and negative otherwise. Patch-text pairs $<\mathbf{z}_{i},\mathbf{t}_{k}>$ are positive if $y_{i}=k$. We construct a loss function of two terms operating on the two types of said pairs:
@@
where $N_{y_{i}}$ is the number of patches in the batch that have the same label as $y_{i}$. This loss term operates on the image patches and it maximizes the similarities of the patches of the same category across different images. This term is a generalization of SupCon [Khosla et al.(2020)Khosla, Teterwak, Wang, Sarna, Tian, Isola, Maschinot, Liu, and Krishnan] to image patches with no data augmentation.
A similar loss function has been proposed for image classification in [Aljundi et al.(2023)Aljundi, Patel, Sulc, Chumerin, and Reino] and it was shown that $\ell_{t}$ (3) formulation can be expressed as a smooth approximation to the maximum function of SupCon and Cross Entropy (CE) loss; in our case of SupCon on patch-text pair and CE on the patch feature with the corresponding text feature as a representative class prototype. This approximation is key to allow a smooth optimization of the different similarities while tolerating possible noisy pairs. We optimize the alignment module $\mathcal{M}$ with $\ell_{\text{TSupCon}}$ (Eq.  1), for a fixed number of epochs and provide ablation on different loss functions in the Appendix. Deployment. At test time, we extract $\{\mathbf{x}_{i}\}$ features from the image, using the frozen image encoder. The image patches are forwarded through $\mathcal{M}$, $\mathbf{z}_{i}=\mathcal{M}(\mathbf{x}_{i})$, then we compute for each $z_{i}$ the most similar text feature $\mathbf{t}_{k}$, after which we assign to $\mathbf{z}_{i}$ a label $k$. Text features are precomputed using a frozen text encoder. To provide pixel level predictions we interpolate the similarities to the original image dimension.
@@ 4.1 Experimental Setup
Implementation Details. We consider Vit-L/14 for both CLIP and DINOv2. We train our alignment module with SGD optimizer, cosine annealing scheduler and a batch size of 5 images. We use COCO-Stuff [Caesar et al.(2018)Caesar, Uijlings, and Ferrari] images for pretraining our alignment model, composed of only 118k images. We extract pseudo segmentations of COCO-Stuff vocabulary using the first stage of our pipeline (section  3.2). Next, we use them to train our alignment module for 10 epochs. Since Stage 1.2 generates many more masks per image compared to Stage 1.1, in order to balance the pseudo annotations of these complementary methods, we randomly sample $75\%$ of annotations from stage 1.2 (see Appendix for ablation).
@@
Figure 3: Qualitative evaluation of Stage.1.1. SAM query points generated by our method are shown in green stars. Left shows instances of correct segmentations by Stage.1.1. and Right demonstrates its limitations; small objects, wrongly detected classes (due to ambiguities) and not enough query points to cover all instances. Stage 1.2 alleviates the issue with small objects and incomplete masks since it labels the masks generated accurately by SAM.
Figure 4: Qualitative results of zeroshot segmentation. The first row shows the ground truth labels. The second row shows the results of FMbSeg.
Datasets. COCO Stuff [Caesar et al.(2018)Caesar, Uijlings, and Ferrari] contains 80 things categories and 91 stuff classes. We evaluate on both things only (CO-80) and things + stuff (CS-171) categories. Pascal VOC [Everingham et al.(2010)Everingham, Van Gool, Williams, Winn, and Zisserman] contains 20 foreground classes and everything else is labeled as background. To isolate the effect of the background class we evaluate on both (PV-20) excluding the background class and (PV-21) with the background class. For PV-21, we add stuff classes from CS-171 as background categories. Pascal Context [Mottaghi et al.(2014)Mottaghi, Chen, Liu, Cho, Lee, Fidler, Urtasun, and Yuille] (PC-59) contains 59 classes of objects and stuff. CityScapes  [Cordts et al.(2016)Cordts, Omran, Ramos, Rehfeld, Enzweiler, Benenson, Franke, Roth, and Schiele] (City) contains 30 categories from street view images. ADE20K [Zhou et al.(2019)Zhou, Zhao, Puig, Xiao, Fidler, Barriuso, and Torralba] (ADE) contains 150 categories from various indoor and outdoor scenes.
@@
GroupVit [Xu et al.(2022a)Xu, De Mello, Liu, Byeon, Breuel, Kautz, and Wang] 41M +Labels 50.4 79.7 18.7 27.5 15.3 11.1 9.2
Mask CLIP [Zhou et al.(2022)Zhou, Loy, and Dai] - 38.8 74.9 23.6 20.6 16.4 12.6 9.8
ReCo [Shin et al.(2022)Shin, Xie, and Albanie] - 25.1 57.7 19.9 31.6 14.8 21.1 11.2
TCL [Cha et al.(2023)Cha, Mun, and Roh] 15M + Labels. 55.0 83.2 33.9 31.6 22.4 24.0 17.1
CLIP-S4 [He et al.(2023)He, Jamonnak, Gou, and Ren] 0.12M - - 33.6 - 22.1 - -
SCCLIP [Wang et al.(2023a)Wang, Mei, and Yuille] - 59.1 - 30.4 - 22.4 - -
CLIP-DIY [Wysoczańska et al.(2024)Wysoczańska, Ramamonjisoa, Trzciński, and Siméoni] - 59.0 - 30.4 - - - -
CaR [Sun et al.(2023)Sun, Li, Torr, Gu, and Li] - 67.6 - 30.5 36.6 - -
SAM-CLIP [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari] 41M 60.6 - 29.2 31.5 - 17.1
FMbSeg 0.12M 67.73 85.65 42.72 57.63 29.88 28.37 16.25
@@
Table 1: Semantic Segmentation performance on various datasets. Best method marked in bold and second best in gray, our method is better or on par with SOTA methods.
@@
Compared Methods. We compare with methods that target the alignment of image features with language for open vocabulary semantic segmentation. We divide methods according to the annotation and training they require. Image Level Annotation: GroupVit [Xu et al.(2022a)Xu, De Mello, Liu, Byeon, Breuel, Kautz, and Wang] and TCL [Cha et al.(2023)Cha, Mun, and Roh]. Annotation-free methods: We consider Mask CLIP [Zhou et al.(2022)Zhou, Loy, and Dai] (the best performing variant), CLIP-S4 [He et al.(2023)He, Jamonnak, Gou, and Ren] and SAM-CLIP [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari]. Training free methods: CLIP-DIY [Wysoczańska et al.(2024)Wysoczańska, Ramamonjisoa, Trzciński, and Siméoni], SCCLIP [Wang et al.(2023a)Wang, Mei, and Yuille], CaR [Sun et al.(2023)Sun, Li, Torr, Gu, and Li] and ReCo [Shin et al.(2022)Shin, Xie, and Albanie].
@@
Metrics. We consider the widely adopted Mean Intersection over Union (mIoU). We follow the evaluation protocol of TCL [Cha et al.(2023)Cha, Mun, and Roh]. We don’t apply any post-refinement and only one standard template {a photo of a “$class$”} for text descriptions is considered unlike other approaches using 80 different templates during evaluation.
4.2 Annotation-Free Semantic Segmentation Results

### v3

@@ 3.2 Stage 1: Object detection & masks generation
3.2.1 Stage 1.1: Querying SAM with CLIP
High-resolution feature extraction We (over-)sample each image into a high resolution one and divide it into $C$ crops in a sliding window fashion. We then process each crop separately with CLIP and rearrange the patches returned from CLIP (section 3.1) for all crops to construct the features for the full image. This guarantees precise and high quality feature map $f_{i}$ for each image. Besides, for each crop $c$, we extract a classification token $\mathbf{cls}_{c}$. We refer to the Appendix for details and visualizations.
Defining the set of concerned categories. To detect classes present in an image our method computes the similarity between the classification token for each crop $\mathbf{cls}_{c}$ and a set of text features $\{\mathbf{t}_{k}\}$ corresponding to descriptions extracted from vocabulary $V$ (section 3.1). All possible labels in a given dataset are considered as the vocabulary (e.g. a set of 171 classes for COCO-Stuff).
Object presence detection. Each crop $c$ is classified with an object category when the object’s text feature $\textbf{t}_{k}$ has the highest similarity to the classification token $\mathbf{cls}_{c}$ among all descriptions extracted from vocabulary $V$. An object category is considered as present in the image if it has been assigned to more than a predefined number of crops (set to 1 in our experiments).
Pseudo mask generation. We compute a similarity matrix between the full image feature map $f_{i}$ and the text features for the detected categories (Fig. 2). For each category $k$ we select 5 patches with the highest similarity as query points. We feed these points along with the original image to SAM and select the mask with the highest confidence $m_{i}^{k}$.
3.2.2 Stage 1.2: SAM masks classification
Stage 1.1 may ignore small objects or generate partial masks for an image (Fig. 3) due to sub-optimal query points. Thus we perform a complementary pseudo label generation mechanism to further boost the performance.
@@
Mask labelling. Given the generated high resolution feature map $f_{i}$ and the detected categories in Stage.1, to classify the generated masks we compute mean feature corresponding to the area covered by each mask and compute its similarity to the text features of the detected categories in the image. The class with the highest similarity is selected as the pseudo label for the corresponding mask $m_{i}^{k}$.
Given the high-resolution feature map $f_{i}$ and the detected categories from Stage 1, we classify the generated masks by computing the mean feature for each mask’s area. We then compare this mean feature to the text features of the detected categories in the image. The class with the highest similarity $class_{k}$is chosen as the pseudo label for the corresponding mask $m_{i}^{k}$.
@@ 3.3 Stage 2: Lightweight semantic segmentation
Stage 1 extracts segmentation masks from a dataset with a predefined vocabulary, but these masks can be noisy. Additionally, querying two foundation models can be inefficient under low compute constraints. To address this, we propose using the annotations generated in Stage 1 to align any off-the-shelf pretrained vision encoder with text semantics, with no human supervision.
@@
Alignment module. We use the generated masks in 3.2 as pseudo labels to train a simple alignment module that maps image patch features to text embeddings. This mapping grounds any pretrained vision encoder with language for dense prediction tasks. To avoid biases in supervised models, we focus on self-supervised pretrained models, specifically DINOv2 [26], which is trained fully self-supervised without text alignment. To handle noisy pseudo annotations, we rely on: 1) frozen pretrained text features as anchors, 2) already discriminative image patch features, and 3) a uniquely designed loss function that is robust to noise.
Pseudo label assignment. Let $\mathbf{D}$ be a dataset of unlabelled images $\mathbf{D}=\{im_{i}\}^{M}_{i=1}$. Using the first stage of the pipeline, we extract object masks for each image $im_{i}$ along with their assigned text features $\{m_{i}^{s},t_{s}\}$. The output of the image encoder (e.g. DINOv2) is $\mathbf{x}_{i}=[\mathbf{x}^{1}_{i},\dots,\mathbf{x}^{p}_{i},\dots,\mathbf{x}^{N}_{i}]$ where $N$ is the number of patches, ignoring the $\mathbf{cls}$ token. Using the generated masks and their associated detected categories we assign to each patch $\mathbf{x}^{p}_{i}$ a pseudo label $y^{p}_{i}\in\{1,..,K\}$ where $K$ is the number of detected categories in the dataset $\mathbf{D}$. From the text encoder we extract $K$ text features $T=\{\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{K}\}$.
@@
Training the alignment module. The pretrained image encoder remains frozen and we optimize a small alignment module $\mathcal{M}$ to map the patch representations $\{\mathbf{x}^{p}_{i}\}$ to the CLIP text embedding space $z^{p}_{i}=\mathcal{M}(\mathbf{x}^{p}_{i})\in\mathbb{R}^{D}$. Note that the method is agnostic to the specific encoder used. As we strive for simplicity, we design our alignment module as one transformer block with multi-head self-attention layer. The self-attention layer allows each patch to attend to other patches in the image. Since we aim for an open-set semantic segmentation, cross-attention with text features [2, 47] is not used as it would require a joint processing of the image features and a closed set of text features at test time.
@@
For notation clarity, we drop the image index and consider only an across-batch patch index $i;i\in\{1,\dots,N*B\}$ where N is the number of patches in an image and $B$ is the batch size. CLIP [18] contrasts the similarity of text features with image class tokens using a cross entropy loss, where a one to one correspondence exists between an image class token and its text features. In our case, we have many image features $\mathbf{z}_{i}=\mathcal{M}(\mathbf{x}_{i})$ extracted from patches of many images and few text features corresponding to detected categories in $\mathbf{D}$. Upon early experiments with CLIP loss, we found it not scalable to image patches and exhibiting poor convergence.
@@
Inspired by supervised contrastive loss SupCon [20] that operates on positive and negative pairs, we construct two types of pairs: pairs of image patches (patch-patch pairs: $<\mathbf{z}_{i},\mathbf{z}_{j}>$) and pairs of image patches and text features (patch-text pairs: $<\mathbf{z}_{i},\mathbf{t}_{k}>$). A patch-patch pair $<\mathbf{z}_{i},\mathbf{z}_{j}>$ is considered positive if the patches belong to the same category class $y_{i}=y_{j}$ and negative otherwise. Patch-text pairs $<\mathbf{z}_{i},\mathbf{t}_{k}>$ are positive if $y_{i}=k$. We construct a loss function of two terms operating on the two types of said pairs:
@@
where $N_{y_{i}}$ is the number of patches in the batch that have the same label as $y_{i}$. This loss term operates on the image patches and it maximizes the similarities of the patches of the same category across different images. This term is a generalization of SupCon [20] to image patches with no data augmentation.
A similar loss function has been proposed for image classification in [3] and it was shown that $\ell_{t}$ (3) formulation can be expressed as a smooth approximation to the maximum function of SupCon and Cross Entropy (CE) loss; in our case of SupCon on patch-text pair and CE on the patch feature with the corresponding text feature as a representative class prototype. This approximation is key to allow a smooth optimization of the different similarities while tolerating possible noisy pairs. We optimize the alignment module $\mathcal{M}$ with $\ell_{\text{TSupCon}}$ (Eq.  1), for a fixed number of epochs and ablate different loss functions in Section 4.4.3.
Deployment. At test time, we extract $\{\mathbf{x}_{i}\}$ features from the image, using the frozen image encoder. The image patches are forwarded through $\mathcal{M}$, $\mathbf{z}_{i}=\mathcal{M}(\mathbf{x}_{i})$, then we compute for each $z_{i}$ the most similar text feature $\mathbf{t}_{k}$, after which we assign to $\mathbf{z}_{i}$ a label $k$. Text features are precomputed using a frozen text encoder.
Pixelwise Segmentation. Since our alignment module works at patch-level, to generate pixel level predictions, we interpolate the similarities to the original image dimension, we call this model FMbSeg-Stage 2 (base). For more refined and accurate segmentation, we leverage SAM. We label the automatically generated masks by SAM using our alignment module, based on the majority vote of the classified patches within each mask. As SAM might miss certain regions depending on the hyperparameters set for mask quality and size, we overlay the interpolated segmentation with the labeled masks for a complete result. We call this model FMbSeg-Stage 2 (refined). Note that any off-the-shelf segmentation model, such as Efficient-SAM [35] or Fast SAM [43], can be used.
@@ 4.1 Experimental Setup
Implementation Details. We consider Vit-L/14 for both CLIP and DINOv2. We train our alignment module with SGD optimizer, cosine annealing scheduler and a batch size of 5 images. We use COCO dataset’s [6] unlabelled images for pretraining our alignment model. Note that other training-based zeroshot semantic segmentation methods use either COCO [8, 18, 9] or a much larger dataset with labels [36, 33] to optimize their model (table 1). We extract pseudo segmentations of COCO-Stuff vocabulary using the first stage of our pipeline (section  3.2). Next, we use them to train our alignment module for 10 epochs only. Our results are generated in a much more constrained setting using only COCO-stuff images and class list without any groundtruth annotations.
@@
Figure 3: Qualitative evaluation of Stage 1.1. SAM query points generated by our method are shown in green stars. Left shows instances of correct segmentations by Stage 1.1. and Right demonstrates its limitations; small objects, wrongly detected classes (due to ambiguities) and not enough query points to cover all instances. Stage 1.2 alleviates the issue with small objects and incomplete masks since it labels all the masks generated accurately by SAM.
Figure 4: Qualitative results of zeroshot segmentation. The first row shows the ground truth labels. The second row shows the results of FMbSeg-Stage 2 (refined).
Datasets. COCO-Stuff [6] contains 80 things categories and 91 stuff classes. We evaluate on both things only (CO-80) and things + stuff (CS-171) categories. Pascal VOC [15] contains 20 foreground classes and everything else is labeled as background. To isolate the effect of the background class we evaluate on both (PV-20) excluding the background class and (PV-21) with the background class. For PV-21, we add stuff classes from CS-171 as background categories in the evaluation of our method where these categories are mapped to the main background class. Pascal Context [25] (PC-59) contains 59 classes of objects and stuff. CityScapes  [11] (City) contains 30 categories from street view images. ADE20K [45] (ADE) contains 150 categories from various indoor and outdoor scenes. This dataset has the least overlap in terms of categories with our training dataset COCO and depicts the zeroshot performance of our method.
@@
GroupVit [36] 41M +Labels 50.4 79.7 18.7 27.5 15.3 11.1 9.2
Mask CLIP [46] - 38.8 74.9 23.6 20.6 16.4 12.6 9.8
ReCo [30] - 25.1 57.7 19.9 31.6 14.8 21.1 11.2
TCL [8] 15M + Labels* 55.0 83.2 33.9 31.6 22.4 24.0 17.1
ZeroSeg [9] 3.4M* - 37.3 19.7 - 17.8 - -
CLIP-S4 [18] 0.12M* - - 33.6 - 22.1 - -
SCCLIP [32] - 59.1 - 30.4 - 22.4 - -
CLIP-DIY [34] - 59.0 - 30.4 - - - -
CaR [31] - 67.6 - 30.5 36.6 - -
SAM-CLIP [33] 41M 60.6 - 29.2 - 31.5 - 17.1
FMbSeg- Stage 2 (base) 0.12M (Stage 1 Annotations)* 67.73 85.65 42.72 57.63 29.88 28.37 16.25
FMbSeg- Stage 2 (refined) 0.12M (Stage 1 Annotations)* 71.02 87.03 44.34 58.20 30.65 30.55 16.60
@@
Table 1: Semantic Segmentation performance on various datasets. Best method underlined, best and second best marked in bold, our method is better or on par with SOTA methods. Methods using COCO dataset as part of their training are marked by *.
@@
Compared Methods. We compare with methods that target the alignment of image features with language for open vocabulary semantic segmentation. We divide methods according to the annotation and training they require. Image Level Annotation: GroupVit [36] and TCL [8]. Annotation-free methods: We consider ZeroSeg [9], Mask CLIP [46] (the best performing variant), CLIP-S4 [18] and SAM-CLIP [33]. Training free methods: CLIP-DIY [34], SCCLIP [32], CaR [31] and ReCo [30].
@@
Metrics. We consider the widely adopted Mean Intersection over Union (mIoU). We follow the evaluation protocol of TCL [8]. We don’t apply any post-refinement and only one standard template {a photo of a “$class$”} for text descriptions is considered unlike other approaches using 80 different templates during evaluation.
4.2 Annotation-Free Semantic Segmentation